Minnie’s Mineral Springs is a single-price monopoly. Columns 1 and 2 of the table set out the market demand schedule for Minnie’s water and columns 2 and 3 set out Minnie’s total cost schedule.
1. Calculate Minnie’s marginal revenue schedule and draw a graph of the market demand curve and Minnie’s marginal revenue curve. Explain why Minnie’s marginal revenue is less than the price.

Price
(dollars per bottle) Quantity demanded
(bottles per hour)
Total cost
(dollars per hour)
10 0 1
8 1 3
6 2 7
4 3 13
2 4 21
0 5 31
2. At what price is Minnie’s total revenue maximized and over what range of prices is the demand for water elastic? Why will Minnie not produce a quantity at which the market demand is inelastic?
3. Calculate Minnie’s profit-maximizing output and price and economic profit.
1
Expert's answer
2019-04-08T09:56:50-0400

P Qd TC TR MR

10 0 1 0 -

8 1 3 8 8

6 2 7 12 4

4 3 13 12 0

2 4 21 8 -4

0 5 31  0 -8

1. Minnie’s marginal revenue is less than the price because it gets less revenue for previous units as well (it has to reduce price to the same amount for all units).

2. TR is maximized, when P = 5. Demand is elastic, when price is between 5 and 10. Minnie will not produce a quantity at which the market demand is inelastic, because MR < MC in this range.

3. Minnie’s profit-maximizing output is at MR = MC, so Q = 2 units and P = 6.

TP = TR - TC = 12 - 7 = 5.